OpenStudy (anonymous):

Somebody please help meee!!! Suppose that U^238 has a half life of 4.5 billion years, decaying (through a series of relatively short lived intermediate atoms) to Pb^206. In a certain mineral sample there are .31 times as many Pb^206 atoms as there are of U^238. If one assumes that the mineral deposit contained no Pb^206 when it was formed and that no lead or uranium have been added to or escaped from the sample (except through the natural decay process) how old is the sample?

OpenStudy (jamesj):

*bookmark

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

JAMES CAN YOU HELP <3

OpenStudy (jamesj):

Right. Having a half life of time T means that after time t, a material will be \[ d(t) = 2^{-t/T} \] of what it was originally. For example at time t= 0, it should be all itself, so \[ d(0) = 2^0 = 1.\] At time t= T, it should have half decayed, so it's half of what it was, \[ d(T) = 2^{-T/T} = 2^{-1} = \frac{1}{2} \] After time t=2T, it should be half of half of the original substance, i.e., 1/4. And that's what our formula gives us, as \[ d(2T) = 2^{-2T/T} = 2^{-2} = \frac{1}{4}. \] So are you convinced this formula d(t) makes sense?

OpenStudy (anonymous):

yes the formula makes sense, i just dont understand how i'm suppose to use it with lead and uranium together ..

OpenStudy (jamesj):

Ok. In this problem, what is the half time, T?

OpenStudy (jamesj):

It's right there at the beginning of your problem: "Suppose that U^238 has a half life of 4.5 billion years, decaying (through a series of relatively short lived intermediate atoms) to Pb^206."

OpenStudy (jamesj):

Hence T = ... what? Talk to me.

OpenStudy (anonymous):

ok so t = 4.5 billion years right?

OpenStudy (jamesj):

The half life T = 4,500,000,000 years, yes. What we want to find is the time t. So let's be careful to differentiate between little t and capital T.

OpenStudy (anonymous):

ok, i'm listening.

OpenStudy (jamesj):

Now, let U denote the amount of Uranium and let Pb denote the amount of iron. What do we know about the ratio: \[ \frac{Pb}{U} \] What is it equal to?

OpenStudy (anonymous):

.31/1

OpenStudy (jamesj):

Right. Now, in terms of U and Pb, what fraction of the original U is still Uranium?

OpenStudy (anonymous):

uhm, after the half life none of it.

OpenStudy (jamesj):

No. The total amount of material now at time t is U + Pb. Therefore what fraction of the total is Uranium?

OpenStudy (jamesj):

If there C cats and D dogs, what fraction of all the animals are cats?

OpenStudy (anonymous):

uhm the ratio multiplied by t right?

OpenStudy (jamesj):

\[ \frac{C}{C+D} \] Therefore what fraction of the total of U + Pb is Uranium?

OpenStudy (anonymous):

okay so itd be C / C+D

OpenStudy (anonymous):

okay so U / U + Pb

OpenStudy (jamesj):

Right, and \[ \frac{U}{U + Pb} = \frac{1}{1 + Pb/U} \] Now, what's that equal to?

OpenStudy (anonymous):

honestly, i wish i could tell you

OpenStudy (jamesj):

We just calculated Pb/U above

OpenStudy (anonymous):

Oh! .31/1

OpenStudy (jamesj):

Pb/U = 0.31 Therefore \[ \frac{U}{U+Pb} = \frac{1}{1+Pb/U} = \frac{1}{1.31} \] Yes?

OpenStudy (anonymous):

ok ya that makes sense.. sorry i'm such a retard btw aha

OpenStudy (jamesj):

Now, what we have is that \[ d(t) = \frac{U}{U + Pb} \] You have a formula for the left-hand side (LHS) and now you know the value of the RHS. What formula do you now have then for t?

OpenStudy (anonymous):

do we have to differentiate?

OpenStudy (jamesj):

Nope. \[ d(t) = 2^{-t/T} \] remember.

OpenStudy (anonymous):

ok so that is equal to what you said above, whats the different between the big t and the little t?

OpenStudy (jamesj):

T = half life t = elapsed time

OpenStudy (anonymous):

because one of them is 4.5 billion years right? and then we just have to solve for the other one? but i dont understand what we sub in for the different elements then

OpenStudy (jamesj):

We have after a time t, \( d(t) \) of the Uranium is still Uranium. We need to solve for t. Now, another expression for the amount of material that is still Uranium is \[ \frac{U}{U + Pb} \] because Pb is what the U decays into. Therefore if we know the value of \( \frac{U}{U + Pb} \) we can find the time t that has passed in order that \[ d(t) = \frac{U}{U + Pb} \]

OpenStudy (anonymous):

and we do know that, so we can solve for t?

OpenStudy (jamesj):

Now that last equation is equivalent to \[ 2^{-t/T} = \frac{1}{1.31} \] You know the value of the half life T. Now solve for elapsed time t.

OpenStudy (anonymous):

ok, when solving should i put the whole 4.5 billion years in or just keep it at 4.5?

OpenStudy (jamesj):

You can use whatever units you like. Time units of billion years is perfectly good.

OpenStudy (anonymous):

so then if i solve for t as 4.5 that should be fine as long as my answer is in billion years?

OpenStudy (jamesj):

T = 4.5. NOT, NOT t. You're solving for t.

OpenStudy (anonymous):

ok right. can i ask one more question? how would i go about solving for the variable i know that if its e with an exponent you just ln it, but i'm not sure if its an actual number.

OpenStudy (jamesj):

\[ 2^{-t/T} = 1/1.31 \] take the log base 2 of both sides and you have \[ \frac{-t}{T} \log_2 2 = \log_2 (1/1.31) \]

OpenStudy (anonymous):

ah ok thank you so much!!!

OpenStudy (jamesj):

What answer do you get for t?

OpenStudy (anonymous):

i got 1.75

OpenStudy (anonymous):

right or wrong?

OpenStudy (jamesj):

Me too. t = 1.75 billion years.

OpenStudy (anonymous):

oh yay i did it right!! ok thank you so much!

OpenStudy (anonymous):

i know i've already asked you a billion questions because your so smart, but do you know how to change basis of transformation matricies?

OpenStudy (jamesj):

Let T be your original matrix with respect to a basis E = {e1, e2, ... , en}. Let F = {f1, f2, ..., fn} be another basis and U the matrix that transforms E to F. Then if you want T in basis F, then you want to use \[ UTU^{-1} \] I think the logic for this is clear. \( U^{-1} \) takes a vector in basis F, puts it in terms of the basis of E. \( T \) then operates on it. Then \( U \) transforms that back into the basis F.

OpenStudy (anonymous):

ok thanks! i'm trying to get something from its basis in F, back to the standard basis!